OpenStudy (anonymous):

Parallel and Perpendicular Lines:Find (a) the slope <-->Xy(Suppose to be under) and (b) the slope of any line perpendicular to <-->XY(Suppose to be under. The Question is: X(-3,4), Y(1,-2)

OpenStudy (anonymous):

slope = (change in y)/(change in x) = (-2-4)/(1--3) = -6/4 = -3/2 slope of perpendicular = - reciprocal of above slope = 2/3

OpenStudy (anonymous):

Slope: -3/2, Perpendicular: 2/3?